A 450 kg satellite is in a circular orbit at an altitude of 525 km above the Earth's surface. Because of air friction, the satellite eventually falls to the Earth's surface, where it hits the ground with a speed of 2.00 km/s. How much energy was transformed to internal energy by means of friction?
Answer #1

For "orbit", centripetal acceleration = gravitational acceleration, or

v^2/r = GM/r^2

v^2 = GM/r

and KE = 1/2*m*v^2

= GMm/2R

and GPE = -GMm/R


so Total Energy = -GMm/2r



so initially total energy

= -6.67*10^-11 * 450*5.98*10^24 / 2*(450+6400)*10^3


TE = -1.310*10^10 J


leter GPE will be

= 6.67*10^-11 * 450*5.98*10^24 / (6400)*10^3

= 2.804*10^10


KE = 1/2*m*v^2 = 1/2 * 450*(2000)^2 = 9*10^8 J


so by conservation of energy


Initail energy = Final energy

-1.310*10^10 = -2.804*10^10 + 9*10^8 + lost energy

lost energy = 1.403*10^10 J
Answer #4

Me = 5.98*10^24 kg
Re = 6370 km = 6.37*10^6 m
h = 525 km = 0.525*10^6 m


intial potential energy U1 = -G*Me*m/(Re+h)

intial kinetic energy k1 = 0.5*m*V^2

= 0.5*m*(G*Me/(Re+h))


initial total mechanical energy, TE1 = U1 + K1

= -0.5*G*Me*m/(Re+h)

= -0.5*6.67*10^-11*5.98*10^24*450/((6.37+0.525)*10^6)

= -1.3*10^10 J

fial potential energy U2 = -G*Me*m/Re

= -6.67*10^-11*5.98*10^24*450/6.37*10^6

= -2.818*10^10 J

fial kinetic energy k1 = 0.5*m*V^2

= 0.5*450*(2000)^2

= 0.09*10^10 J


final total mechanical energy, TE2 = U2 + K2

= -2.728*10^10 J


workdone by friction = TE2 - TE1

= -2.728*10^10 - (-1.3*10^10)

= -1.428*10^10 J